# T2VEval: Benchmark Dataset and Objective Evaluation Method for T2V-generated Videos

### v5

@@ V-C2 Ablation Experiment of Realness Branch Pre-train Strategy
V-C3 Ablation Experiment of The Effectiveness of Each Branch and Training Strategy
In this section, we performed ablation experiments to evaluate the effectiveness of each branch and the divide-and-conquer training strategy employed. Specifically, we removed each of the three branches of the model and conducted fine-tuning tests on the T2VEval-Bench dataset. The dataset partitioning method remains consistent with the previous experiments. The results, presented in Table IV, demonstrate that removing any branch leads to varying degrees of performance degradation, confirming the importance and effectiveness of the three dimensions assessed in the T2V quality evaluation task. Notably, removing the technical quality branch has the smallest impact on model performance, suggesting that when evaluating text-to-video generation systems, focusing solely on technical quality may prove insufficient. A comprehensive assessment framework should additionally prioritize ”content quality evaluation”, encompassing two critical dimensions we proposed: (1) visual realness and (2) text-video alignment.
Additionally, we evaluated the impact of the divide-and-conquer training strategy we proposed by comparing it with a single-stage training approach where the entire model is trained at once using the overall impression MOS scores (denoted as T2VEval-Once). The results, shown in Table III, indicate that our divide-and-conquer training strategy outperforms the single-stage approach. This finding highlights that the divide-and-conquer strategy effectively facilitates cooperation between branches, leading to improved overall performance.
VI Conclusion
In conclusion, this paper focuses on the benchmark dataset and quality evaluation for text-to-video generation. To this end, we constructed a multi-dimensional text-to-video quality evaluation dataset, T2VEval-Bench, which comprises 1,783 videos generated by 13 advanced T2V models. Through subjective experiment, we collected the mean opinion scores on four dimensions: the overall impression, the text-video consistency, the realness, and the technical quality. Based on T2VEval-Bench, we developed a multi-branch fusion framework for T2V quality evaluation called T2VEval, which extracts video features from three key perspectives: text-video alignment, technical quality, and realness. These features are fused using an attention-based mechanism, and the final scoring prediction is performed by using a large language model (LLM). To enhance training efficiency, a divide-and-conquer strategy is employed, enabling collaborative training across the branches. Experimental results demonstrate that T2VEval can effectively evaluates the quality of T2V generated videos. In future work, we plan to expand the dataset by including additional models and keeping up-to-date with model version updates. Dataset and codes will be open-sourced upon completion of the follow-up work.
{refcontext}
[…]

### v6

@@ 5.3.2 Ablation Experiment of Realness Branch Pre-train Strategy
5.3.3 Ablation Experiment of The Effectiveness of Each Branch and Training Strategy
In this section, we performed ablation experiments to evaluate the effectiveness of each branch and the divide-and-conquer training strategy employed. Specifically, we removed each of the three branches of the model and conducted fine-tuning tests on the T2VEval-Bench dataset. The dataset partitioning method remains consistent with the previous experiments. The results, presented in Table 5, demonstrate that removing any branch leads to varying degrees of performance degradation, confirming the importance and effectiveness of the three dimensions assessed in the T2V quality evaluation task. Notably, removing the technical quality branch has the smallest impact on model performance, suggesting that when evaluating text-to-video generation systems, focusing solely on technical quality may prove insufficient. A comprehensive assessment framework should additionally prioritize "content quality evaluation", encompassing two critical dimensions we proposed: (1) visual realness and (2) text-video alignment.
Additionally, we evaluated the impact of the divide-and-conquer training strategy we proposed by comparing it with a single-stage training approach where the entire model is trained at once using the overall impression MOS scores (denoted as T2VEval-Once). The results, shown in Table 4, indicate that our divide-and-conquer training strategy outperforms the single-stage approach. This finding highlights that the divide-and-conquer strategy effectively facilitates cooperation between branches, leading to improved overall performance.
6 Conclusion
Suffering from the infringement of the patent right of cartoon characters, we aim to propose an effective scheme to reduce the illegal use of cartoon characters. Considering the demands in patent examination and other application scenarios, we built our cartoon character dataset CCDaS with sufficient annotations, which both annotated the face and the whole-body of the cartoon characters. As for our proposed CCD algorithm MP-YOLO, we not only maintained the real-time detection speed and small parameter size of YOLO, but achieved a high detection result for our CCD task. As a result, our approach is competitive with the state-of-the-art and can be effectively used in the patent examination process or other application scenarios. In the future, we would like to explore other CCD methods and continuously improve the accuracy of CCD.
Acknowledgments
This paper is supported by "the China National Science and Technology Major Project (2022ZD0116306)".